Tano issues bonds with a par value of $92,000 on January 1, 2017. The bonds annual contract rate is 10%, and interest is pailaho Issues bonds with a par Value OT $92,000 on January 1, 2017. The bonds annual contact rate is 10%, and Interest is paldTano issues bonds with a par value of $92,000 on January 1, 2017. The bonds annual contract rate is 10%, and interest is pai
Answer #1
1
Discount 4520 =92000-87480
2
Total interest expense over life of bonds
6 payments of $ 4600 27600
Par value at maturity 92000
Total repaid 119600
Less: Amount borrowed 87480
Total bond interest expense 32120
3
Semiannual Interest period end Unamortized Discount Carrying value
01/01/2017 4520 87480
06/30/2017 3767 88233
12/31/2017 3014 88986
06/30/2018 2261 89739
12/31/2018 1508 90492
06/30/2019 755 91245
12/31/2019 0 92000
Workings:
Semiannual Interest 4600 =92000*10%*6/12
Semiannual Discount amortized 753 =4520/6